### v1

# Improving the Reasoning of Multi-Image Grounding in MLLMs via Reinforcement Learning

Abstract
Recently, Multimodal Large Language Models (MLLMs) excel at visual grounding in single-image scenarios with textual references. However, their performance degrades when handling real-world applications involving complex multi-image compositions and multimodal instructions, which reveals limitations in cross-image reasoning and generalization. To address these challenges, we adopt a Reinforcement Learning (RL) based post-training strategy to improve the reasoning performance of MLLMs in multi-image grounding tasks. Our approach begins with synthesizing high-quality chain-of-thought (CoT) data for cold-start initialization, followed by supervised fine-tuning (SFT) using low-rank adaptation (LoRA). The cold-start training stage enables the model to identify correct solutions. Subsequently, we perform rejection sampling using the merged SFT model to curate high-quality RL data and leverage rule-based RL to guide the model toward optimal reasoning paths. Extensive experimental results demonstrate the effectiveness of our approach, achieving +9.04% improvements on MIG-Bench and +4.98% improvements on several out-of-domain reasoning grounding benchmarks over the SFT baseline. Furthermore, our approach exhibits strong generalization in multi-image perception, with gains of +3.1% and +2.4% over the base model on subsets of the BLINK and MMIU benchmarks, respectively.
1 Introduction
Figure 1: Top: Examples of our multi-image reasoning grounding. Bottom: The comparison of Qwen2.5-VL-7B model, SFT model, and our reasoning model. Our method achieves the best performance among all benchmarks.
The traditional visual grounding task focuses on localizing target regions within a single image using simple natural language descriptions. Recent research leveraging large language models (LLMs) has significantly improved performance on this task by utilizing LLMs’ powerful language comprehension capabilities. However, real-world applications often require sophisticated multi-image grounding with complex instructions, including cross-image reasoning. As illustrated in Figure 1, practical scenarios often require identifying the region that best matches the semantics of the query image. Such tasks demand advanced multi-modal reasoning and comprehensive understanding across multiple visual contexts.
Recently, Migician li2025migician introduces a multi-image grounding dataset with the MIG-Bench benchmark. By employing a two-stage supervised fine-tuning (SFT) pipeline on their dataset, they have achieved impressive results, demonstrating the potential of MLLMs for multi-image grounding tasks. However, their SFT-based approach primarily “memorize” chu2025sft understanding patterns and instruction-following behaviors in multi-image scenarios, rather than developing robust reasoning capabilities. This limitation constrains the model’s ability to generalize to novel scenarios and real-world applications.
Inspired by the recent success in RL-based post-training frameworks for Large Reasoning Models (LRM)  guo2025deepseek ; team2025kimi ; yang2025qwen3 , we explore the potential of the RL paradigm to improve the reasoning abilities of MLLMs in multi-image grounding scenarios. Our preliminary attempts applied RL directly to the state-of-the-art open-source MLLM, Qwen2.5-VL-7B bai2025qwen2 . However, we found that the model often failed to generate correct responses during the RL training stage, which we attribute to two key limitations of the base model: (1) inadequate capacity for processing complex image queries, and (2) insufficient ability to handle multi-image contexts. It suggests that directly employing RL paradigm on the base model is ineffective for exploring reasoning paths in such challenging scenarios.
To overcome these limitations, we synthesize high-quality chain-of-thought (CoT) data for cold-start initialization using large-scale MLLMs, Qwen2.5-VL-72B. This data serves for supervised fine-tuning (SFT) with low-rank adaptation (LoRA), enabling the model to develop robust multi-image comprehension and reasoning capabilities. Then, we remove low-quality data through rejection sampling according to predictions from the merged cold-start SFT model. Finally, we employ a rule-based RL paradigm based on Group Relative Policy Optimization (GRPO) shao2024deepseekmath to guide the model toward discovering correct reasoning paths while improving generalization performance.
We conduct comprehensive experiments on several benchmarks to demonstrate the effectiveness of our training pipeline. On the multi-image grounding benchmark MIG-Bench li2025migician , our model achieves improvements of +3.46% (in-domain) and +12.76% (out-of-domain) over the SFT baseline. For zero-shot grounding evaluation, our approach outperforms the SFT baseline by +6.23% on Ref-L4 chen2025revisiting , +4.79% on LISA-Grounding lai2024lisa , and +2.83% on ReVOS-Grounding yan2024visa . Furthermore, for the multi-image understanding benchmarks, our approach exhibits strong generalization capabilities, with improvements of +3.1% on BLINK fu2024blink and +2.4% on MMIU meng2024mmiu compared to the base Qwen2.5-VL-7B model.
2 Related Works
2.1 MLLM-based Referring Grounding.
Referring grounding, also known as referring expression comprehension (REC), involves identifying a specific object within an image based on a given referring natural language expression. With the rapid advancement of multimodal large language models (MLLMs) liu2023visual ; wang2024qwen2 ; li2024llava ; jiang2024mantis ; li2024llavainterleave ; bai2025qwen2 ; zhu2025internvl3 , significant progress has been made in extending their capabilities to address more challenging tasks, including fine-grained image understanding, object detection zhan2024griffon ; jiao2024lumen , and visual grounding lai2024lisa ; yan2024visa ; rasheed2024glamm ; li2024groundinggpt . Although these MLLMs have achieved strong performance on standard benchmarks such as RefCOCO/+/g mao2016generation ; yu2016modeling , their visual grounding capabilities remain largely confined to single-image contexts and face substantial difficulties in capturing cross-image relationships. Recent efforts, such as Migician li2025migician , introduce multi-image grounding tasks and further propose a two-stage supervised fine-tuning strategy, achieving notable results in these tasks. However, such approaches still exhibit insufficient reasoning abilities and limited generalization in multi-image and out-of-domain settings. In this paper, we explore the potential of the RL paradigm to improve the reasoning and comprehension abilities of MLLMs in multi-image grounding scenarios.
2.2 Reinforcement Learning-based Reasoning Model
Recent advancements in improving the reasoning ability of large language models jaech2024openai ; xu2024llava ; el2025competitive ; guo2025deepseek ; team2025kimi have been made through Chain-of-Thought (CoT) training or reinforcement learning-based post-training. Previous methods, such as Reinforcement Learning from Human Feedback (RLHF) ouyang2022training and Direct Preference Optimization (DPO) rafailov2023direct , have explored aligning models with human preferences, achieving remarkable success in enhancing reasoning and instruction-following capabilities. Recently, DeepSeek-R1 guo2025deepseek employed the Group Relative Policy Optimization (GRPO) shao2024deepseekmath algorithm as a post-training method without the need for critic models, demonstrating the potential of large-scale reinforcement learning. Inspired by the success of DeepSeek-R1, a series of subsequent works rapidly adopted this approach to enhance multi-modal reasoning across various domains, including mathematical problem solving yang2025r1 ; peng2025skywork ; huang2025vision ; zhang2025r1 ; deng2025openvlthinker , video understanding feng2025video ; zhang2025tinyllava ; liao2025improved ; wang2025time , and visual perception tasks shen2025vlm ; liu2025seg ; deng2025boosting ; liu2025visual ; zhan2025vision ; yu2025perception ; tan2025reason . Among the most closely related works to ours, VLM-R1 shen2025vlm , Visual-RFT liu2025visual , Vision-R1 zhan2025vision , Perception-R1 yu2025perception , most focus exclusively on single-image visual grounding or object detection. However, these approaches lack the capability for multi-image perception and reasoning. In this paper, we follow the DeepSeek-R1 paradigm and extend it to the multi-image grounding tasks, while also achieving strong performance on single-image grounding and multi-image understanding tasks.
3 Method
3.1 Basic Settings
We begin by formally defining the multi-image grounding task. Formally, given a natural language description $t$, a query image $Q$, and several target images $\{I_{1},I_{2},...,I_{m}\}$, the model M is required to generate a bounding box $G$ that indicates the region satisfying the semantic and contextual constraints specified by $t$ and $Q$.
Unlike conventional grounding tasks that rely on fixed instruction templates, multi-image grounding involves dynamic and flexible instructions, which demand significantly stronger instruction comprehension and reasoning capabilities from MLLMs. To address the challenge, we introduce a reinforcement learning training strategy to enhance the model’s reasoning capabilities, inspired by recent breakthrough improvements in LRMs. We employ Qwen2.5-VL-7B as the base model due to its excellent multi-modal comprehension performance. Our RL paradigm is similar to DeepSeek-R1 guo2025deepseek , consisting of cold-start CoT-SFT initialization, rejection sampling, and rule-based RL training. We will provide detailed descriptions of our dataset construction strategy and training paradigm in the following subsections.
3.2 Cold Start Initialization
As analyzed in DeepSeek-R1 guo2025deepseek , directly applying reinforcement learning (RL) from scratch may cause the model to enter an unstable state. Specifically, the base MLLM model struggles to handle complex multi-image contexts and accurately localize target regions within multi-image scenarios. To mitigate these issues, it is crucial to pre-equip the model with task-specific reasoning capabilities before RL fine-tuning. We achieve this by constructing a high-quality CoT cold-start dataset, which provides structured reasoning guidance to bootstrap the model’s initial performance.
Task Subsets Number Source
Mulit-image Grounding Common Object 22k MGrounding-630k
Referring Grounding 24k
Region Locating 5k
Static Difference 3k
Single-image Grounding - 1k Refcoco/g/+
1k ODINW, V3DET
Table 1: Details of constructed CoT cold-start dataset.
3.2.1 Chain-of-Thought Reasoning Data Collection
We construct our CoT dataset primarily based on MGrounding-630k dataset li2025migician . To maintain the model’s capability in single-image grounding scenarios, additionally introduce a small amount of single-image grounding data, including RefCOCO/+/g mao2016generation ; yu2016modeling , as well as data from other sources, such as ODINW li2022grounded and V3Det wang2023v3det .
We utilize the advanced large-scale MLLM Qwen2.5-VL-72B to generate CoT reasoning data. Given a question, along with images and a fixed prompt similar to DeepSeek-R1, the model generates CoT rationales processes in the following format:
  • {Question} First output the thinking process in <think> </think> tags and then output the final answer in <answer> </answer> tags. Output the bounding box coordinates in JSON format.
To improve data quality, we conduct an additional data filtering strategy. For each sample, we prompt the model to produce four distinct responses. Then, We estimate the difficulty of each sample based on the Intersection over Union (IoU) accuracy among these four responses, retaining only samples for which the accuracy is 1 to ensure the reliability of the data. After filtering, we obtain 56k high-reliability CoT cold-start samples, with detailed statistics presented in Table 1.
3.2.2 Supervised Fine-tuning
Leveraging our CoT cold-start dataset, we perform fine-tuning of the Qwen2.5-VL model to develop its multi-image comprehension and reasoning capabilities. During this training phase, we implement strong supervision via next-token prediction across the entire generation process, effectively incorporating knowledge at both the reasoning trace level and the final bounding box prediction.
We recognize that conventional full-parameter fine-tuning risks locking the model into rigid reasoning patterns, which could compromise both its generalization capacity and conversational fluency. To address this, we employ the LoRA fine-tuning approach, which maintains the model’s foundational knowledge while significantly enhancing its perception and reasoning performance in multi-image grounding tasks. Then we merge the LoRA parameters with the original base model, yielding our stage-1 CoT-SFT model.
3.3 Reinforcement Learning
In the second stage, we implement GRPO shao2024deepseekmath , a rule-based reinforcement learning training paradigm, to further enhance the model’s reasoning capabilities. The model samples multiple candidate responses for each sample and are encouraged to reinforce responses achieving higher task-level rewards during GRPO training stage. To be specific, our model generates a group of $G$ complete responses $\{o_{1},o_{2},...,o_{G}\}$ with current parameter group $\pi_{\theta}$. Each response contains a full CoT reasoning process and a final bounding box prediction. For each $o_{i}$, we compute a scalar reward $r_{i}$, and normalize these rewards to estimate group-relative advantages $A_{i}$, formally as:
$A_{i}=\frac{r_{i}-mean(\{r_{j}\}_{j=1}^{N})}{std(\{r_{j}\}_{j=1}^{N})}$ (1)
Then, the training objective is given as follows:
$\mathcal{J}_{GRPO}(\theta)=\frac{1}{N}\sum_{i=1}^{N}(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}A_{i}-\beta·\mathcal{KL}(\pi_{\theta}(o_{i}|q)|\pi_{ref}(o_{i}|q))$ (2)
where $\pi_{\theta_{old}}$ and $\pi_{ref}$ are the parameter groups of old model without latest updating and the fixed stage-1 CoT-SFT model. The parameter $\epsilon$ is used to control the clipping range, and $\beta$ is the KL penalty coefficient.
Rejection Sampling
To ensure efficient training, we implement a rejection sampling strategy based on our stage-1 CoT-SFT model. Specifically, we perform multiple sampling on the RL training data and remove samples where the model’s predictions are all correct or wrong. This approach is adopted because samples that the model answers either completely incorrectly or completely correctly do not contribute effectively to the relative advantage estimation in the GRPO algorithm. Retaining only samples where the model provides partially correct answers enables more efficient model optimization. After rejection sampling, we collect 174k data for stage-2 training.
To optimize the training efficiency of our reinforcement learning framework, we employ a rejection sampling strategy using our stage-1 CoT-SFT model. Specifically, we generate multiple predictions for each sample in the RL training dataset. Those samples where all predictions are correct or wrong are removed, as they contribute poorly to the relative advantage estimation in GRPO. On the other hand, partially correct samples offer optimal learning signals for model improvement. After applying this rigorous filtering process, we obtain a high-quality dataset of 174,000 samples for stage-2 RL training.
Reward Function
We employ two reward functions for the multi-image grounding task: accuracy reward $r_{acc}$ and format reward $r_{format}$.
Accuracy Reward. Given the ground truth bounding box $B$ and the predicted bounding box $\hat{B}$, we calculate the $IoU(B,\hat{B})$ as our accuracy reward, where $IoU$ denotes the Intersection over Union metric.
Format Reward. We adopt a format reward similar to that used in DeepSeek-R1, where the model output must follow the format: “<think>…</think><answer>…</answer>”. The reward score is set to 1 only when the output adheres to the required format; otherwise, the score is 0. Additionally, we impose constraints on the bounding box to ensure it is output in the correct JSON format.
The total reward is computed as a weighted sum of the accuracy and format rewards:
$r=\lambda_{acc}r_{acc}+\lambda_{format}r_{format}$ (3)
where $\lambda_{acc}$ and $\lambda_{format}$ denote the weights assigned to the accuracy reward and the format reward, respectively.
4 Experiments
Models Spontaneous Grounding Referential Grounding AVG
Difference Similarity Visual Reference Textual Visual+Textual
Static Robust Common OT MV Region Refer GG Reason Co-Re
Human Performance
Human 99.50* 97.87 98.00* 100.00 96.88 100.00* 98.99 91.06* 92.08 97.44 97.18
70B-Scale MLLMs
LLaVA-OV-72B 13.26 5.34 26.84 12.91 7.64 2.14 17.83 21.60 11.88 8.55 13.65
InternVL2-76B 15.91 10.64 36.40 30.73 20.83 5.74 46.46 41.28 32.67 26.50 26.72
InternVL3-78B 10.04 9.57 24.12 27.08 14.58 10.44 50.51 38.08 45.54 17.09 24.71
Qwen2-VL-72B 46.12 46.81 64.46 26.73 22.57 18.62 33.33 62.53 50.50 17.09 38.88
Qwen2.5-VL-72B 43.75 46.81 69.98 34.32 29.17 8.31 62.63 59.92 66.34 41.03 46.23
7B-Scale MLLMs
Mantis 1.52 0.00 3.31 12.18 2.08 1.00 1.01 10.02 0.00 0.85 3.20
LLaVA-OV-7B 6.06 3.19 3.43 0.18 1.04 1.08 9.09 15.43 6.93 0.85 4.73
Minicpm2.6 14.58 2.13 14.34 9.82 6.25 1.75 11.11 10.02 2.97 2.56 7.55
mPLUG-Owl3 18.56 6.38 34.93 8.55 7.64 2.41 7.07 22.85 9.09 5.98 12.35
InternVL2-8B 6.92 7.45 25.49 20.73 9.72 3.49 28.28 30.26 17.82 9.40 15.96
InternVL3-8B 23.67 14.89 47.99 14.84 6.94 12.13 7.07 34.87 16.83 2.56 18.18
Qwen2-VL-7B 27.84 38.30 19.36 20.73 11.81 25.95 23.23 58.52 48.51 11.97 28.62
Qwen2.5-VL-7B 28.03 5.32 21.81 15.23 5.56 4.07 13.13 32.26 3.96 2.56 13.19
+ SFT 53.22 44.68 82.11 26.59 35.07 39.32 79.80 59.72 53.47 23.93 49.79
+ CoT-SFT 47.35 45.74 76.59 46.14 34.03 28.93 80.81 63.93 66.34 34.19 52.41
Ours-GRPO 51.52 43.62 82.48 63.41 38.54 51.45 82.83 69.54 67.33 37.61 58.83
Table 2: Performance comparison of different models on MIG-Bench. OT, MV, GG, and Co-Re respectively mean object tracking, multi-view grounding, group grounding, and correspondence. For values marked with *, we randomly sample 20% testing examples for human evaluation on the corresponding task. The best and second-best results are marked in bold and underline, respectively.
Models In-Domain Out-of-Domain AVG
Qwen2.5-VL-72B 46.17 46.27 46.23
Qwen2.5-VL-7B 16.76 10.82 13.19
+ SFT 63.61 40.58 49.79
+ CoT-SFT 58.42 48.40 52.41
Ours 67.07 53.34 58.83
Table 3: In-domain vs Out-of-domain performance comparison on MIG-Bench li2025migician . The in-domain subsets include static difference, common object recognition, region localization, and referential grounding, while the out-of-domain subsets comprise robust difference, object tracking, multi-view grounding, group grounding, and correspondence.
Model Single image Grounding Multi-image Grounding AVG
Ref-L4(val) Ref-L4(test) LISA(val) LISA(test) ReVOS
Qwen2.5-VL-72B 74.85 76.71 53.57 46.85 48.78 60.15
Qwen2.5-VL-7B 58.57 62.76 56.63 53.54 30.07 52.31
+ SFT 68.92 69.14 55.10 54.18 46.33 58.73
Ours 74.78 75.74 62.76 56.11 49.16 63.71
Table 4: Zero-shot evaluation on several reasoning grounding benchmarks: Ref-L4 chen2025revisiting , LISA-Grounding lai2024lisa , and ReVOS yan2024visa . The best and second-best results are marked in bold and underline, respectively.
In this section, we first describe the implementation details of our approach in Section 4.1: benchmarks, baselines, and evaluation metrics. Subsequently, we present the comparison results in Section 4.2.
4.1 Implementation Details
Benchmarks
We evaluate our method on the multi-image grounding benchmark MIG-Bench li2025migician , object detection dataset ODINW li2022grounded , single-image grounding datasets: Ref-L4 chen2025revisiting , LISA-Grounding lai2024lisa , and the video grounding dataset ReVOS Grounding yan2024visa . ODINW comprises 13 distinct real-world settings featuring rare object categories, assessing the model’s object perception and inference ability in practical scenarios. Moreover, to further evaluate out-of-domain generalization ability of our model, we individually employ five non-overlapping subsets from ODINW: Oxford Pets, Pascal VOC, Wildfire Smoke, Open Poetry Vision, and Vector. Ref-L4 is a large-scale referring expression benchmark with diverse object categories, instance scales, and long textual descriptions, making it ideal for assessing reasoning capabilities of MLLMs. For ReVOS, we uniformly sample six frames and task the model with grounding one frame, deriving ground-truth bounding boxes from segmentation masks originally annotated for reasoning tasks.
To demonstrate the generalizability of our RL training strategies across multi-image understanding tasks, we evaluate our approach on two additional benchmarks: BLINK fu2024blink and MMIU meng2024mmiu . For BLINK, we assess performance on seven relevant subsets. For MMIU, we evaluate on four relevant types: low-level semantic relations, high-level semantic relationships (objective), high-level semantic relationships (subjective), and two-dimensional spatial relationships.
Training Configurations
In the cold start CoT-SFT stage, we employ a learning rate of 1e-4 with cosine decay scheduling and an accumulated total batch size of 32. The maximum sequence length for generation is set to 1,024 tokens. During this phase, we freeze the vision encoder and MLP projector, updating only the LLM parameters via LoRA. For the subsequent RL phase, we only train the LLM. We set a learning rate of 5e-5 with 8 rollout samples per input, a batch size of 2, and gradient accumulation steps of 4. The reward weights $\lambda_{acc}$ and $\lambda_{format}$ are set to 1.0 and 0.5, respectively. The KL penalty coefficient $\beta$ is set to 0.001, the sampling temperature to 0.7, and the output length remains 1024 tokens. All experiments are conducted with H800-80G GPUs.
Method OxfordPets PascalVOC WildfireSmoke OpenPoetryVision Vector AVG
Qwen2.5-VL-72B 19.36 90.56 38.78 16.85 61.11 45.33
Qwen2.5-VL-7B 18.52 90.56 26.53 15.59 83.33 46.91
+ SFT 19.22 91.26 34.69 15.05 88.89 49.82
Ours 20.61 91.73 45.58 18.37 94.44 54.15
Table 5: Out-of-domain object detection evaluation on ODNIW li2022grounded dataset. We individually employ five non-overlapping subsets: Oxford Pets, Pascal VOC, Wildfire Smoke, Open Poetry Vision, and Vector. The best result is indicated in bold.
Model Counting Forensic IQ Jigsaw Multi-view Similarity Vis. Corr. AVG
Closed-source Model
GPT4o 49.17 79.55 31.33 55.33 59.40 72.59 75.00 60.34
Open-source Model
Qwen2.5-VL-7B 67.50 45.45 27.33 63.33 51.88 82.96 68.02 58.07
+ SFT-63k 63.33 32.58 26.67 72.67 48.12 90.37 62.21 56.56
Ours-new 70.00 47.73 32.00 74.67 49.62 84.44 69.77 61.18
Table 6: Performance on Multi-image Understanding Benchmark BLINK fu2024blink . Forensic, IQ, Multi-view, Similarity, and Vis. Corr. denote Forensic detection, IQ test, Multi-view reasoning, Visual similarity, and Vision Correspondence, respectively. The best and second-best results are marked in bold and underline, respectively. Results for GPT4o  gpt4o are borrowed from BLINK.
Baselines
We select several open-source MLLMs for comparison, such as Qwen2-VL wang2024qwen2 , Qwen2.5-VL bai2025qwen2 , InternVL2 chen2024internvl , InternVL3 zhu2025internvl3 , LLaVA-OneVision li2024llava , MiniCPM2.6 yao2024minicpm , and mPLUG-Owl3 ye2024mplug . To further verify the effectiveness of the RL paradigm, we compare our method with the SFT baseline, which is trained on the same dataset as our stage-1 CoT-SFT with a learning rate of 2e-6 and an accumulated batch size of 64.
Evaluation Metrics
We employ the standard Acc@0.5 metric for the grounding task, which considers a prediction correct if its Intersection over Union (IoU) with the ground truth exceeds 0.5.
4.2 Comparison Results
MIG-Bench Comparison
As shown in Table 2, we present the comparison results in the multi-image grounding benchmark MIG-Bench li2025migician . Our method achieves state-of-the-art performance overall. Specifically, it obtains the best results across all seven subsets of referential grounding. Compared to the base model (Qwen2.5-VL-7B), our approach achieves a 45.64% improvement, and it outperforms the second-best model (Qwen2.5-VL-72B) by 12.60%, while using significantly fewer parameters.
Moreover, we conduct a comprehensive comparison of various post-training strategies on both in-domain and out-of-domain tasks, shown in Table 3. The results reveal that conventional SFT training significantly improves in-domain performance but provides limited benefits for out-of-domain generalization. CoT-SFT training strategy enhances out-of-domain generalization to a certain extent but slightly compromises in-domain accuracy. Conversely, the RL training strategy achieves superior performance, obtaining the highest scores in both in-domain and out-of-domain settings, with a notable 53.34% on out-of-domain tasks. The results demonstrate the strong generalization capability of our approach across both conventional and novel task domains.
Model High-level Obj. High-level Sub. Low-level AVG
VGR STS VCor SC FC ER CR MIC VQA FD
Closed-source Model
GPT4o 61.3 84.0 72.6 41.5 42.6 46.5 67.8 93.5 79.2 88.8 67.8
Open-source Model
Mantis 65.2 65.5 30.9 28.1 24.3 31.8 61.5 71.0 59.8 57.0 49.5
Llava-interleave 49.8 54.5 25.1 26.4 23.2 24.8 29.5 57.0 48.8 26.3 36.5
Qwen2.5-VL-7B 68.0 74.0 44.6 33.6 25.1 31.0 57.0 80.0 82.8 72.0 56.8
+ SFT 67.0 80.0 40.3 28.6 26.1 32.5 69.2 87.5 81.2 65.1 57.7
Ours 66.8 72.0 51.1 30.4 27.0 39.0 69.5 87.0 82.2 69.3 59.4
Table 7: Performance on Multi-image Understanding Benchmark MMIU meng2024mmiu . High-level Obj. and High-level Sub. are represented as subtypes of high-level semantic relationship (objective) and high-level semantic relationship (subjective). The term "Low-level" denotes the subtype of low-level semantic relations. Some results are borrowed from the original paper. The best and second-best results among open-source models are marked in bold and underline, respectively.
Model Two-dimensional spatial relationship AVG
RPM JPS ICSC IQASC ITRSC ISTE HE PT SOT
Closed-source Model
GPT4o 12.5 23.5 58.5 88.0 17.5 35.0 29.5 55.0 56.0 41.7
Open-source Model
Mantis 13.5 27.5 48.0 66.0 13.0 23.5 23.0 76.0 50.0 37.8
Llava-interleave 15.0 17.5 47.5 42.0 14.0 22.5 21.5 79.0 60.5 35.5
Qwen2.5-VL-7B 13.0 26.5 55.5 75.5 7.0 29.0 23.5 58.5 55.5 37.8
+ SFT 12.5 30.5 54.0 73.0 19.5 33.5 29.5 43.0 52.5 38.6
Ours 15.0 24.5 68.0 83.5 19.0 32.5 22.0 50.0 49.5 40.4
Table 8: Performance on Multi-image Understanding Benchmark MMIU meng2024mmiu for two-dimensional spatial relationship subtype. Some results are borrowed from the original paper. The best and second-best results among open-source models are marked in bold and underline, respectively.
Zero-shot Evaluation
To further highlight the robust generalization capability of our method, we conduct zero-shot evaluation on multiple datasets spanning different tasks: the single-image grounding dataset Ref-L4 chen2025revisiting , LISA lai2024lisa , and the video grounding dataset ReVOS yan2024visa . As illustrated in Table 4, our method consistently outperforms all baselines across multiple benchmarks. Although the SFT training strategy achieves slight overall improvements, it exhibits a severe performance drop on the LISA benchmark. It sugguests that the SFT training method lacks strong generalization, relying on memorizing answers instead of learning underlying patterns. On the other hand, our RL training strategy, which is initialized with a small amount of data, achieves a substantial improvement of 4.98% over the SFT baseline and even surpasses the performance of Qwen2.5-VL-72B by 3.56%.
Out-of-domain Object Detection
The evaluation results on out-of-domain object detection task are presented in Table 5. As described in Section 4.1, we evaluate on five non-overlapping datasets from ODNIW, comprising 6,314 samples. Our full model achieves the highest average performance (54.15%) across all five datasets, surpassing the base model Qwen2.5-VL-7B by a significant improvement of 7.24%. Notably, it also outperforms both the SFT baseline and the largest scale base model, Qwen2.5-VL-72B. It demonstrate the robustness and strong generalization capability of our approach in handling out-of-domain object detection tasks.
Multi-image Understanding Evaluation
Our method not only advances multi-image reasoning and grounding capabilities but also exhibits exceptional generalization in the multi-image understanding task. For comprehensive evaluation, we conduct extensive experiments on the BLINK fu2024blink benchmark, with seven representative subsets: Counting, Forensic Detection, IQ Test, Jigsaw, Multi-view Reasoning, Visual Similarity, and Visual Correspondence. As shown in Table 6, our RL-based method achieves state-of-the-art performance, outperforming both the base model Qwen2.5-VL-7B bai2025qwen2 and the SFT baseline. What’s more, our method surpasses even the closed-source GPT-4o model gpt4o , establishing new performance standards in multi-image understanding. Notably, although our post-training dataset includes only grounding tasks, the model demonstrates remarkable performance on multi-image understanding tasks. It reveals that the reasoning process during reinforcement learning can enhance the multi-image comprehension capabilities of the model.
For comprehensive evaluation on the MMIU meng2024mmiu benchmark, we assess performance across semantic relationships and two-dimensional spatial relationships, as shown in Table 7 and Table 8, respectively. We compare our method with several closed-source models: GPT4o gpt4o , open-source models: Mantis jiang2024mantis , Llava-interleave li2024llavainterleave , and base model Qwen2.5-VL-7B bai2025qwen2 . We achieve the best performance among open-source models across all types.
Ablation Setting Spontaneous Grounding Referential Grounding AVG
CS RS RL Static Robust Common OT MV Region Refer GG Reason Co-Re
Qwen2.5-VL-7B - - - 28.03 5.32 21.81 15.23 5.56 0.54 13.13 32.26 3.96 2.56 14.05
[…]
Table 9: Ablations on MIG-Bench. CS and RS are represented as cold-start and rejection sampling. RL denotes Reinforcement learning.
4.3 Ablation Studies
To demonstrate the effectiveness of our approach, we perform ablation studies on MIG-Bench li2025migician , shown in Table 9.
Effectiveness of Cold Start on GPRO
As discussed in Section 3.2, we directly apply GRPO-based reinforcement learning following the DeepSeek-R1-Zero framework guo2025deepseek . However, this attempt yields only minor performance gains, primarily due to the base model’s constraints in multi-image contexts and complex queries. These limitations underscore the critical importance of implementing a CoT-based cold-start strategy to establish a strong initialization for subsequent reward-based RL learning.
Effectiveness of Rejection Sampling
Another key component of our approach is the data filtering via rejection sampling. To assess its impact, we conduct an ablation study by training the stage-1 CoT-SFT model directly with the GRPO algorithm, removing the rejection sampling process. As illustrated in the fourth row of Table 9, our full model—with rejection sampling enabled—achieves a improvement of 3.27%. It shows the importance of data quality in reinforcement learning training and demonstrates the effectiveness of our rejection sampling procedure.
5 Conclusion
In this paper, we aim to enhance the reasoning and understanding capabilities of MLLMs for real-world multi-image grounding applications. To this end, we propose a post-training strategy incorporating cold-start initialization to establish correct reasoning pathways, complemented by rule-based reinforcement learning to further strengthen the model’s reasoning abilities. Experimental results demonstrate that our approach significantly outperforms both the base model and the baseline SFT method across multiple multi-image grounding and understanding benchmarks. We hope this work will stimulate further research in the multi-image reasoning and grounding community.