### v2

# Improving the Reasoning of Multi-Image Grounding in MLLMs via Reinforcement Learning

Abstract
Recently, Multimodal Large Language Models (MLLMs) excel at visual grounding in single-image scenarios with textual references. However, their performance degrades when handling real-world applications that involve complex multi-image compositions and multi-modal instructions, revealing limitations in cross-image reasoning and generalization. To address these challenges, we adopt a Reinforcement Learning (RL) based post-training strategy to improve the reasoning of MLLMs in multi-image grounding tasks. Our approach begins with synthesizing high-quality chain-of-thought (CoT) data for cold-start initialization, followed by supervised fine-tuning (SFT) using low-rank adaptation (LoRA). The cold-start training stage enables the model to identify correct solutions. Subsequently, we perform rejection sampling using the merged SFT model to curate high-quality RL data and leverage rule-based RL to guide the model toward optimal reasoning paths. Extensive experimental results demonstrate the effectiveness of our approach, yielding improvements of +9.04% on MIG-Bench, +6.37% on MC-Bench, and +4.98% on several out-of-domain reasoning grounding benchmarks compared to the SFT baseline. Furthermore, our method exhibits strong generalization in multi-image perception, with gains of +3.1% and +2.4% over the base model on BLINK and MMIU benchmarks, respectively.
The traditional visual grounding task focuses on localizing target regions within a single image using simple natural language descriptions. Recent research leveraging large language models (LLMs) has significantly improved performance on this task through their powerful language comprehension capabilities. However, real-world applications often require sophisticated multi-image grounding with complex instructions, including cross-image reasoning. As illustrated in Figure 1 (a), practical scenarios often require identifying the region that best matches the semantics of the query image. Such tasks demand advanced multi-modal reasoning and comprehensive multi-image understanding.
Recently, Migician li2025migician introduces a multi-image grounding dataset and employs a two-stage supervised fine-tuning (SFT) pipeline on their dataset. However, their SFT approach primarily “memorize” chu2025sft understanding patterns and instruction-following behaviors in multi-image scenarios, rather than developing reasoning capabilities. This limitation constrains the model’s generalization to novel scenarios and real-world applications.
1 Introduction
Figure 1: (a): Example of our multi-image reasoning grounding results. (b): The comparison of the Qwen2.5-VL-7B model, its SFT variant, and our reasoning model. Our method achieves the best performance among all benchmarks.
Inspired by the recent success in RL-based post-training frameworks for Large Reasoning Models (LRM)  guo2025deepseek ; team2025kimi ; yang2025qwen3 , we explore the potential of the RL paradigm to improve MLLMs’ reasoning abilities in multi-image grounding scenarios. Our preliminary attempts applied RL directly to the superior open-source MLLM, Qwen2.5-VL-7B bai2025qwen2 . However, we found that the model often failed to generate correct responses during RL training, due to the limited capacity of the base model to process complex image queries and multi-image contexts.
To address these limitations, we introduce a two-stage training framework, consisting of cold-start CoT-SFT initialization and rule-based reinforcement learning. Firstly, we synthesize high-quality chain-of-thought (CoT) data using a large-scale MLLM, Qwen2.5-VL-72B, and perform supervised fine-tuning with LoRA for cold-start initialization, enabling the model to acquire basic multi-image reasoning capabilities. Next, we adopt a rule-based RL paradigm built on Group Relative Policy Optimization (GRPO) shao2024deepseekmath to guide the model toward discovering correct reasoning paths and improving generalization. However, we observe that a significant portion of data is low-quality, preventing the model from generating response candidates with sufficient relative advantage gaps, thereby providing limited supervisory signals for RL. To improve training efficiency, we apply rejection sampling, according to predictions from the merged cold-start SFT model, to filter out such uninformative samples before the RL stage.
We conduct comprehensive experiments on several benchmarks to demonstrate the effectiveness of our training pipeline. As shown in Figure 1 (b), our model achieves improvements of +3.46% (in-domain) and +12.76% (out-of-domain) over the SFT baseline on the multi-image grounding benchmark MIG-Bench li2025migician . As for the other multi-image grounding benchmark MC-Bench xu2024mc , our model demonstrates a +6.37% performance gain compared to the SFT baseline. For zero-shot grounding evaluation, our approach outperforms the SFT baseline by +4.79% on LISA-Grounding lai2024lisa , +6.23% on Ref-L4 chen2025revisiting , and +2.83% on ReVOS-Grounding yan2024visa . Furthermore, for the multi-image understanding benchmarks, our approach demonstrates strong generalization capabilities, with improvements of +3.1% on BLINK fu2024blink and +2.4% on MMIU meng2024mmiu compared to the base Qwen2.5-VL-7B model.
2 Related Works
2.1 MLLM-based Referring Grounding.
Referring grounding, also known as referring expression comprehension (REC), involves identifying a specific object within an image based on a given natural language referring expression. With the rapid advancement of multimodal large language models (MLLMs) liu2023visual ; wang2024qwen2 ; li2024llava ; jiang2024mantis ; li2024llavainterleave ; bai2025qwen2 ; zhu2025internvl3 , significant progress has been made in extending their capabilities to address more challenging tasks, including fine-grained image understanding wang2024advancing , object detection zhan2024griffon ; jiao2024lumen , and visual grounding lai2024lisa ; yan2024visa ; rasheed2024glamm ; li2024groundinggpt . Although these MLLMs have achieved strong performance on standard benchmarks such as RefCOCO/+/g mao2016generation ; yu2016modeling , their visual grounding capabilities remain largely confined to single-image contexts and face substantial difficulties in capturing cross-image relationships. Recent efforts, such as Migician li2025migician , introduce multi-image grounding tasks and further propose a two-stage SFT strategy, achieving notable results in these tasks. However, such approaches still exhibit insufficient reasoning abilities and limited generalization in multi-image and out-of-domain settings. In this paper, we explore the potential of the RL paradigm to improve the reasoning and comprehension abilities of MLLMs in multi-image grounding scenarios.
2.2 RL-based Reasoning Model
Recent advancements in improving the reasoning ability of large language models jaech2024openai ; xu2024llava ; el2025competitive ; guo2025deepseek ; team2025kimi have been made through CoT-based or RL-based post-training. Previous methods, such as Reinforcement Learning from Human Feedback (RLHF) ouyang2022training and Direct Preference Optimization (DPO) rafailov2023direct , have explored aligning models with human preferences, achieving remarkable success in enhancing reasoning and instruction-following capabilities. Recently, DeepSeek-R1 guo2025deepseek employes the GRPO shao2024deepseekmath algorithm as a post-training method without the need for critic models, demonstrating the potential of large-scale reinforcement learning. Inspired by the success of DeepSeek-R1, a series of subsequent works rapidly adopt this approach to enhance multi-modal reasoning across various domains, including mathematical problem solving yang2025r1 ; peng2025skywork ; huang2025vision ; zhang2025r1 ; deng2025openvlthinker , video understanding feng2025video ; zhang2025tinyllava ; liao2025improved ; wang2025time , and visual perception tasks shen2025vlm ; liu2025seg ; deng2025boosting ; liu2025visual ; zhan2025vision ; yu2025perception ; tan2025reason . Among the most closely related works to ours, VLM-R1 shen2025vlm , Visual-RFT liu2025visual , Vision-R1 zhan2025vision , and Perception-R1 yu2025perception , primarily focus on single-image visual grounding or object detection. However, these approaches lack the capability for multi-image perception and reasoning. In this paper, we follow the DeepSeek-R1 paradigm and extend it to the multi-image grounding tasks, while also achieving strong performance on both multi-image grounding and multi-image understanding tasks.
3 Method
3.1 Preliminaries
We begin by formally defining the multi-image grounding task. Given a natural language description $t$, query images $Q$, and a set of $m$ target images $\{I_{1},I_{2},...,I_{m}\}$, the model M is required to localize a region of interest by generating a bounding box $G$ that satisfying the semantic and contextual constraints specified by $t$ and $Q$.
While traditional grounding tasks typically rely on fixed instruction templates, the multi-image grounding task introduces the additional complexity of processing dynamic, context-sensitive instructions. This advancement necessitates MLLMs with substantially enhanced comprehension and reasoning abilities. To bridge this capability gap, we adopt an RL-based training paradigm, leveraging recent advances in LRMs, which have demonstrated significant improvements in complex reasoning tasks. Specifically, we choose Qwen2.5-VL-7B as the base model due to its excellent multi-modal comprehension performance. As shown in Figure 2, our training paradigm consists of two stages: cold-start CoT-SFT initialization and rule-based RL training. We will provide detailed descriptions of each stage in the following subsections.
3.2 Cold-start CoT-SFT Initialization
As analyzed in DeepSeek-R1 guo2025deepseek , directly applying reinforcement learning from scratch may result in unstable convergence. Specifically, the base model struggles to handle complex multi-image contexts and accurately localize target regions in multi-image scenarios. Consequently, it is essential to pre-equip the MLLM with task-specific reasoning capabilities before RL training. To address this, we construct a high-quality CoT cold-start dataset that provides reasoning guidance to bootstrap the model’s initial performance.
Figure 2: The overview of our two-stage training paradigm. Our training paradigm consists of a cold-start CoT-SFT initialization (stage-1) and a rule-based RL training (stage-2).
CoT Reasoning Dataset
Our CoT reasoning dataset is constructed based on MGrounding-630k dataset li2025migician . To maintain the model’s capability in single-image grounding scenarios, we augment our dataset with a small amount of single-image grounding data RefCOCO/+/g mao2016generation ; yu2016modeling , as well as object detection datasets: ODINW li2022grounded and V3Det wang2023v3det . This hybrid approach enables comprehensive training across both multi-image and single-image grounding tasks.
We employ a large-scale MLLM Qwen2.5-VL-72B to generate the CoT reasoning data. Following the paradigm of DeepSeek-R1, our generation process takes a question-image pair accompanied by a fixed instructional prompt as input. The MLLM then produces structured CoT rationales in the following format:
  • {Question} First output the thinking process in <think> </think> tags and then output the final answer in <answer> </answer> tags. Output the bounding box coordinates in JSON format.
To establish high-quality training data, we implement an additional filtering strategy. For each sample, we prompt the Qwen2.5-VL-72B model to generate four distinct responses. Next, we assess sample consistency by computing the Intersection over Union (IoU) accuracy across all responses. To guarantee maximum reliability in our training set, we retain only samples achieving perfect agreement (IoU accuracy = 1.0). This filtering process yields 56k high-confidence CoT cold-start samples, with detailed statistics provided in Table 1.
Task Subsets Number Source
Mulit-image Grounding Common Object 22k MGrounding-630k
Referring Grounding 24k
Region Locating 5k
Static Difference 3k
Single-image Grounding - 1k Refcoco/g/+
1k ODINW, V3DET
Table 1: Details of constructed CoT cold-start dataset.
Cold-start Supervised Fine-tuning
Leveraging our curated CoT dataset, we equip the Qwen2.5-VL-7B model with multi-image comprehension and reasoning capabilities through cold-start initialization. To be specific, we perform strong supervision via next-token prediction across the entire generation process. This stage ensures comprehensive learning of both the reasoning trace and its spatial grounding predictions.
We recognize that full-parameter fine-tuning may constrain the model’s reasoning flexibility, potentially diminishing both its generalization capabilities and the quality of its natural language interactions. To mitigate these limitations, we employ a LoRA fine-tuning approach, which selectively updates only low-rank decomposed matrices. This approach maintains the model’s foundational knowledge while significantly enhancing its perception and reasoning performance in multi-image grounding tasks. After fine-tuning, we merge the learned LoRA modules with the base MLLM, resulting in our Stage-1 CoT-SFT model with improved capabilities.
3.3 Rule-based RL Training
In the second stage, we implement group relative policy optimization (GRPO) shao2024deepseekmath , a rule-based reinforcement learning algorithm designed to improve policy optimization by incorporating group-wise relative constraints on policy updates. To improve training efficiency, we integrate rejection sampling before GRPO, filtering out uninformative samples to accelerate convergence and stabilize learning.
Group Relative Policy Optimization
The MLLM model generates a group of $G$ complete responses $\{o_{1},o_{2},...,o_{G}\}$ with current policy $\pi_{\theta}$. The output response is required to contain a full CoT reasoning process and a final bounding box prediction. For each response $o_{i}$, we compute a scalar reward $r_{i}$, and normalize these rewards to estimate its group-relative advantage $A_{i}$, formally as:
$A_{i}=\frac{r_{i}-mean(\{r_{j}\}_{j=1}^{N})}{std(\{r_{j}\}_{j=1}^{N})}$ (1)
Then, the GRPO training objective can be defined as follows:
$\mathcal{J}_{GRPO}(\theta)=\frac{1}{N}\sum_{i=1}^{N}(\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{old}}(o_{i}|q)}A_{i}-\beta·\mathcal{KL}(\pi_{\theta}(o_{i}|q)|\pi_{ref}(o_{i}|q))$ (2)
where $\pi_{\theta_{old}}$ is the previous policy before update, and $\pi_{ref}$ is the fixed CoT-SFT policy after stage-1 training. The hyperparameters $\epsilon$ and $\beta$ control the policy update clipping range and KL divergence penalty strength, respectively.
Models Referential Grounding Spontaneous Grounding AVG
Visual Reference Textual Visual+Textual Difference Similarity
OT MV Region Refer GG Reason Co-Re Static Robust Common
Human Performance
Human 100.00 96.88 100.00* 98.99 91.06* 92.08 97.44 99.50* 97.87 98.00* 97.18
70B-Scale MLLMs
LLaVA-OV-72B 12.91 7.64 2.14 17.83 21.60 11.88 8.55 13.26 5.34 26.84 13.65
InternVL2-76B 30.73 20.83 5.74 46.46 41.28 32.67 26.50 15.91 10.64 36.40 26.72
InternVL3-78B 27.08 14.58 10.44 50.51 38.08 45.54 17.09 10.04 9.57 24.12 24.71
Qwen2-VL-72B 26.73 22.57 18.62 33.33 62.53 50.50 17.09 46.12 46.81 64.46 38.88
Qwen2.5-VL-72B 34.32 29.17 8.31 62.63 59.92 66.34 41.03 43.75 46.81 69.98 46.23
7B-Scale MLLMs
Mantis 12.18 2.08 1.00 1.01 10.02 0.00 0.85 1.52 0.00 3.31 3.20
LLaVA-OV-7B 0.18 1.04 1.08 9.09 15.43 6.93 0.85 6.06 3.19 3.43 4.73
Minicpm2.6 9.82 6.25 1.75 11.11 10.02 2.97 2.56 14.58 2.13 14.34 7.55
mPLUG-Owl3 8.55 7.64 2.41 7.07 22.85 9.09 5.98 18.56 6.38 34.93 12.35
InternVL2-8B 20.73 9.72 3.49 28.28 30.26 17.82 9.40 6.92 7.45 25.49 15.96
InternVL3-8B 14.84 6.94 12.13 7.07 34.87 16.83 2.56 23.67 14.89 47.99 18.18
Qwen2-VL-7B 20.73 11.81 25.95 23.23 58.52 48.51 11.97 27.84 38.30 19.36 28.62
Qwen2.5-VL-7B 15.23 5.56 4.07 13.13 32.26 3.96 2.56 28.03 5.32 21.81 13.19
+ SFT 26.59 35.07 39.32 79.80 59.72 53.47 23.93 53.22 44.68 82.11 49.79
+ CoT-SFT 46.14 34.03 28.93 80.81 63.93 66.34 34.19 47.35 45.74 76.59 52.41
Ours 63.41 38.54 51.45 82.83 69.54 67.33 37.61 51.52 43.62 82.48 58.83
Table 2: Performance comparison of different models on MIG-Bench. OT, MV, GG, and Co-Re respectively mean object tracking, multi-view grounding, group grounding, and correspondence. For values marked with *, we randomly sample 20% testing examples for human evaluation on the corresponding task. The best and second-best results are marked in bold and underline, respectively.
Rejection Sampling
We implement rejection sampling leveraging our stage-1 CoT-SFT model to enhance the quality of training data. This selective approach involves: (1) generating multiple predictions for each sample in the RL training dataset, and (2) systematically eliminating samples where the model demonstrates either complete correctness (all predictions correct) or failure (all predictions incorrect). The rejection sampling is theoretically grounded in the requirements of the GRPO algorithm, where samples with uniformly correct or incorrect predictions provide negligible gradient information for relative advantage estimation. By retaining only samples with partial correctness, we ensure the training data contains optimal learning signals that effectively drive model optimization. After filtering, we curate a high-quality dataset comprising precisely 174k samples for stage-2 rule-based RL training.
Reward Design
We employ two reward functions for the multi-image grounding task: accuracy reward $r_{acc}$ and format reward $r_{format}$.
Accuracy Reward. Given the ground truth bounding box $B$ and the predicted bounding box $\hat{B}$, we calculate the $IoU(B,\hat{B})$ as our accuracy reward, where $IoU$ denotes the Intersection over Union metric.
Format Reward. We adopt a format reward similar to that used in DeepSeek-R1, where the model output must follow the format: “<think>…</think><answer>…</answer>”. The reward score is set to 1 only when the output adheres to the required format; otherwise, the score is 0. Additionally, we impose constraints on the bounding box to ensure it is output in the correct JSON format.
The total reward is computed as a weighted sum of both the accuracy reward and the format reward:
$r=\lambda_{acc}r_{acc}+\lambda_{format}r_{format}$ (3)
where $\lambda_{acc}$ and $\lambda_{format}$ denote the weights assigned to the accuracy reward and the format reward, respectively.
4 Experiments
Models In-Domain Out-of-Domain AVG
Qwen2.5-VL-72B 46.17 46.27 46.23
Qwen2.5-VL-7B 16.76 10.82 13.19
+ SFT 63.61 40.58 49.79
+ CoT-SFT 58.42 48.40 52.41
Ours 67.07 53.34 58.83
Table 3: In-domain vs Out-of-domain comparison on MIG-Bench li2025migician . The in-domain subsets include static difference, common object recognition, region localization, and referential grounding, while the out-of-domain subsets comprise robust difference, object tracking, multi-view grounding, group grounding, and correspondence. The best result is indicated in bold.
Model Comparison Reasoning Referring $\rm{AP_{50}}$
70B-Scale MLLMs
InternVL2-76B 42.66 35.91 52.78 40.01
InternVL3-78B 45.54 43.11 46.23 43.96
Qwen2.5-VL-72B 71.79 57.95 71.57 64.74
7B-Scale MLLMs
InternVL2-8B 33.90 24.89 34.21 29.28
InternVL3-8B 57.19 37.60 59.05 47.35
Qwen2.5-VL-7B 29.61 20.55 24.23 24.37
+ SFT 60.37 45.02 48.42 52.12
+ CoT-SFT 64.66 45.43 60.89 55.10
Ours 68.93 49.37 59.49 58.49
Table 4: The Comparison on multi-image grounding benchmark MC-Bench xu2024mc . Our method achieve the best performance among 7B-Scale MLLMs, marked in bold.
Model Multi-image Grounding Single image Grounding AVG
ReVOS LISA(val) LISA(test) Ref-L4(val) Ref-L4(test)
Qwen2.5-VL-72B 48.78 53.57 46.85 74.85 76.71 60.15
Qwen2.5-VL-7B 30.07 56.63 53.54 58.57 62.76 52.31
+ SFT 46.33 55.10 54.18 68.92 69.14 58.73
Ours 49.16 62.76 56.11 74.78 75.74 63.71
Table 5: Zero-shot evaluation on several reasoning grounding benchmarks: ReVOS yan2024visa , LISA-Grounding lai2024lisa , and Ref-L4 chen2025revisiting . The best and second-best results are marked in bold and underline, respectively.
Model Counting Forensic IQ Jigsaw Multi-view Similarity Vis. Corr. AVG
GPT4o (Closed-source) 49.17 79.55 31.33 55.33 59.40 72.59 75.00 60.34
Qwen2.5-VL-7B 67.50 45.45 27.33 63.33 51.88 82.96 68.02 58.07
+ SFT 63.33 32.58 26.67 72.67 48.12 90.37 62.21 56.56
Ours 70.00 47.73 32.00 74.67 49.62 84.44 69.77 61.18
Table 6: Performance on Multi-image Understanding Benchmark BLINK fu2024blink . Forensic, IQ, Multi-view, Similarity, and Vis. Corr. denote Forensic detection, IQ test, Multi-view reasoning, Visual similarity, and Vision Correspondence, respectively. The best and second-best results are marked in bold and underline, respectively. Results for GPT4o  gpt4o are borrowed from BLINK.
Method OP PV WS OPV Vec AVG
Qwen2.5-VL-72B 19.36 90.56 38.78 16.85 61.11 45.33
Qwen2.5-VL-7B 18.52 90.56 26.53 15.59 83.33 46.91
+ SFT 19.22 91.26 34.69 15.05 88.89 49.82
Ours 20.61 91.73 45.58 18.37 94.44 54.15
Table 7: Object detection evaluation on ODINW li2022grounded dataset. We individually employ five non-overlapping subsets: OP (Oxford Pets), PV (Pascal VOC), WS (Wildfire Smoke), OPV (Open Poetry Vision), and Vec (Vector). The best result is indicated in bold.
In this section, we first describe the implementation details of our approach in Section 4.1: benchmarks, baselines, and evaluation metrics. Subsequently, we present the comparison results in Section 4.2.
4.1 Implementation Details
Benchmarks
We evaluate our method on the multi-image grounding benchmark MIG-Bench li2025migician , MC-Bench xu2024mc , object detection dataset ODINW li2022grounded , single-image grounding datasets: LISA-Grounding lai2024lisa , Ref-L4 chen2025revisiting , and the video grounding dataset ReVOS Grounding yan2024visa . MIG-Bench is a manually curated benchmark designed to evaluate the multi-image grounding ability of MLLMs, which comprises 4.3k instances and covers 10 distinct tasks. MC-Bench features 2K manually annotated samples, containing three distinct styles of text prompts. To align with our experimental setup, we select samples containing one instance per image for evaluation. ODINW comprises 35 distinct real-world settings featuring rare object categories, assessing the model’s object perception and inference ability in practical scenarios. To further evaluate the generalization ability of our model, we individually employ five non-overlapping subsets from ODINW: Oxford Pets, Pascal VOC, Wildfire Smoke, Open Poetry Vision, and Vector. To align with our setting, we select only those samples that contain a single bounding box per image, resulting in a total of 6,314 samples. Ref-L4 is a large-scale (45k annotations) referring expression benchmark with diverse object categories, instance scales, and long textual descriptions, making it ideal for assessing the reasoning capabilities of MLLMs. For ReVOS, we uniformly sample six frames and task the model with grounding one frame, deriving ground-truth bounding boxes from segmentation masks originally annotated for reasoning tasks.
To demonstrate the generalizability of our RL training strategies across multi-image understanding tasks, we evaluate our approach on two additional benchmarks: BLINK fu2024blink and MMIU meng2024mmiu . For BLINK, we assess performance on seven relevant subsets. For MMIU, we evaluate on four relevant types: low-level semantic relations, high-level semantic relationships (objective), high-level semantic relationships (subjective), and two-dimensional spatial relationships.
Training Configurations
In the cold start CoT-SFT stage, we employ a learning rate of 1e-4 with cosine decay scheduling and an accumulated total batch size of 32. The maximum sequence length for generation is set to 1,024 tokens. During this phase, we freeze the vision encoder and MLP projector, updating only the LLM parameters via LoRA. For the subsequent RL phase, we only train the LLM. We set a learning rate of 5e-5 with 8 rollout samples per input, a batch size of 2, and gradient accumulation steps of 4. The reward weights $\lambda_{acc}$ and $\lambda_{format}$ are set to 1.0 and 0.5, respectively. The KL penalty coefficient $\beta$ is set to 0.001, the sampling temperature to 0.7, and the output length remains 1024 tokens. All experiments are conducted with H800-80G GPUs.
Baselines
We select several open-source MLLMs for comparison, such as Qwen2-VL wang2024qwen2 , Qwen2.5-VL bai2025qwen2 , InternVL2 chen2024internvl , InternVL3 zhu2025internvl3 , LLaVA-OneVision li2024llava , MiniCPM2.6 yao2024minicpm , and mPLUG-Owl3 ye2024mplug . To further verify the effectiveness of the RL training, we compare our method with the SFT baseline, which is trained on the same dataset as our stage-1 CoT-SFT with a learning rate of 2e-6 and an accumulated batch size of 64.
Evaluation Metrics
We employ the standard Acc@0.5 metric for the grounding task, which considers a prediction correct if its Intersection over Union (IoU) with the ground truth exceeds 0.5.
Model High-level Sub. High-level Obj. Low-level Two-dimensional Spatial Relationship Average
ISTE HE PT RPM JPS ICSC IQASC ITRSC SOT VGR STS VCor SC FC ER CR MIC VQA FD Sem Spa
GPT4o(Closed-source) 46.5 67.8 93.5 61.3 84.0 72.6 41.5 42.6 79.2 88.8 12.5 23.5 58.5 88.0 17.5 35.0 29.5 55.0 56.0 67.8 41.7
Mantis 31.8 61.5 71.0 65.2 65.5 30.9 28.1 24.3 59.8 57.0 13.5 27.5 48.0 66.0 13.0 23.5 23.0 76.0 50.0 49.5 37.8
Llava-interleave 24.8 29.5 57.0 49.8 54.5 25.1 26.4 23.2 48.8 26.3 15.0 17.5 47.5 42.0 14.0 22.5 21.5 79.0 60.5 36.5 35.5
Qwen2.5-VL-7B 31.0 57.0 80.0 68.0 74.0 44.6 33.6 25.1 82.8 72.0 13.0 26.5 55.5 75.5 7.0 29.0 23.5 58.5 55.5 56.8 37.8
+ SFT 32.5 69.2 87.5 67.0 80.0 40.3 28.6 26.1 81.2 65.1 12.5 30.5 54.0 73.0 19.5 33.5 29.5 43.0 52.5 57.7 38.6
Ours 39.0 69.5 87.0 66.8 72.0 51.1 30.4 27.0 82.2 69.3 15.0 24.5 68.0 83.5 19.0 32.5 22.0 50.0 49.5 59.4 40.4
Table 8: Performance on Multi-image Understanding Benchmark MMIU meng2024mmiu . High-level Obj. and High-level Sub. are represented as subtypes of high-level semantic relationship (objective) and high-level semantic relationship (subjective). The term "Low-level" denotes the subtype of low-level semantic relations. Sem represents the average score across High-level Object, High-level Subject, and Low-level subtype, while Spa denotes the average score for the Two-dimensional Spatial Relationship subtype. Some results are borrowed from the original paper. The best and second-best results among open-source models are marked in bold and underline, respectively.
Ablation Setting Referential Grounding Spontaneous Grounding AVG
CS RS RL OT MV Region Refer GG Reason Co-Re Static Robust Common
Qwen2.5-VL-7B - - - 15.23 5.56 4.07 13.13 32.26 3.96 2.56 28.03 5.32 21.81 13.19
[…]
Table 9: Ablations on MIG-Bench. CS and RS are represented as cold-start and rejection sampling. RL denotes Reinforcement learning.
Ablation Referential Spontaneous AVG
w/o KL 54.20 57.95 55.33
w. KL, $\beta$=1e-1 51.59 57.15 53.26
w. KL, $\beta$=1e-2 54.96 57.89 54.95
w. KL, $\boldsymbol{\beta}$=1e-3 (Ours) 58.16 59.21 58.47
w. KL, $\beta$=1e-4 56.45 58.86 57.17
w. KL, $\beta$=1e-5 52.29 56.98 53.70
Table 10: Ablations for the weight of KL divergence.
4.2 Comparison Results
Multi-image Grounding Evaluation
As shown in Table 2, we present the comparison results in the multi-image grounding benchmark MIG-Bench li2025migician . Our method achieves state-of-the-art performance overall. Specifically, it obtains the best results across all seven subsets of referential grounding. Compared to the base model (Qwen2.5-VL-7B), our approach achieves a 45.64% improvement, and it outperforms the second-best model (Qwen2.5-VL-72B) by 12.60%, while using significantly fewer parameters. We further evaluate our method on MC-Bench xu2024mc , as shown in Table 4. Our model outperforms all 7B-scale MLLMs, the SFT baseline, and the stage-1 CoT-SFT model. Remarkably, it also surpasses several 70B-scale models, including InternVL2-76B and InternVL3-78B, highlighting its strong reasoning and comprehension capabilities.
Moreover, we conduct a comprehensive comparison of various post-training strategies on both in-domain and out-of-domain tasks, shown in Table 3. The results reveal that conventional SFT training significantly improves in-domain performance but provides limited benefits for out-of-domain generalization. CoT-SFT training strategy enhances out-of-domain generalization to a certain extent but slightly compromises in-domain accuracy. Conversely, our full model achieves superior performance, obtaining the highest scores in both in-domain and out-of-domain settings, with a notable 53.34% on out-of-domain tasks. The results demonstrate the strong generalization capability of our approach across both conventional and novel task domains.
Zero-shot Evaluation
To further highlight the robust generalization capability of our method, we conduct zero-shot evaluation on multiple datasets spanning different tasks: the single-image grounding dataset LISA lai2024lisa , Ref-L4 chen2025revisiting , and the video grounding dataset ReVOS yan2024visa . As illustrated in Table 5, our method consistently outperforms all baselines across multiple benchmarks. Although the SFT training strategy achieves slight overall improvements, it exhibits a severe performance drop on the LISA benchmark. It sugguests that the SFT training method lacks strong generalization, relying on memorizing answers instead of learning underlying patterns. On the other hand, our RL training strategy, which is initialized with a small amount of data, achieves a substantial improvement of 4.98% over the SFT baseline and even surpasses the performance of Qwen2.5-VL-72B by 3.56%.
Object Detection Evaluation
The evaluation results on the object detection dataset ODINW are presented in Table 7. Our full model achieves the highest average performance (54.15%) across all five datasets, surpassing the base model Qwen2.5-VL-7B by a significant improvement of 7.24%. Notably, it also outperforms both the SFT baseline and the largest scale base model, Qwen2.5-VL-72B. It demonstrates the robustness and strong generalization capability of our approach in handling out-of-domain object detection tasks.
Multi-image Understanding Evaluation
Our method not only advances multi-image reasoning and grounding capabilities but also exhibits exceptional generalization in the multi-image understanding task. For comprehensive evaluation, we conduct extensive experiments on the BLINK fu2024blink benchmark, with seven representative subsets: Counting, Forensic Detection, IQ Test, Jigsaw, Multi-view Reasoning, Visual Similarity, and Visual Correspondence. As shown in Table 6, our RL-based method achieves state-of-the-art performance, outperforming both the base model Qwen2.5-VL-7B bai2025qwen2 and the SFT baseline. What’s more, our method surpasses even the closed-source GPT-4o model gpt4o , establishing new performance standards in multi-image understanding. Notably, although our post-training dataset includes only grounding tasks, the model demonstrates remarkable performance on multi-image understanding tasks. It reveals that the reasoning process during reinforcement learning can enhance the multi-image comprehension capabilities of the model.
For comprehensive evaluation on the MMIU meng2024mmiu benchmark, we assess performance across semantic relationships and two-dimensional spatial relationships, as shown in Table 8. We compare our method with several closed-source models: GPT4o gpt4o , open-source models: Mantis jiang2024mantis , Llava-interleave li2024llavainterleave , and base model Qwen2.5-VL-7B bai2025qwen2 . Although our method was not trained on relevant tasks, it achieves the best performance among open-source models across all types.
4.3 Ablation Studies
To demonstrate the effectiveness of our approach, we perform ablation studies on MIG-Bench li2025migician , shown in Table 9.
Effectiveness of Cold Start
As discussed in Section 3.2, we directly apply GRPO-based RL training following the DeepSeek-R1-Zero framework guo2025deepseek , as shown in the second row of Table 9. However, this attempt yields only minor performance gains (+3.74% over the base model), primarily due to the base model’s constraints in handling multi-image contexts and complex queries. These limitations highlight the critical importance of implementing a CoT cold-start strategy to establish a strong initialization for subsequent reward-based RL training. As presented in the third and fourth rows of Table 9, we employ a CoT cold-start initialization using both full-parameter SFT and LoRA-based SFT, respectively. The results indicate that the cold-start strategy substantially improves the model’s performance. Ultimately, we adopt the LoRA-based cold start approach due to its slightly better performance (+0.59% compared to full-parameter SFT).
Effectiveness of Rejection Sampling
Another key component of our approach is data filtering via rejection sampling. To assess its impact, we conduct an ablation study by training the stage-1 CoT-SFT model directly with the GRPO algorithm, removing the rejection sampling process. As illustrated in the fourth row of Table 9, our full model (with rejection sampling enabled) achieves an improvement of 3.27%. It shows the importance of data quality in RL training and demonstrates the effectiveness of our rejection sampling procedure.
Ablations for the weight of KL divergence
Additionally, we conduct an ablation study on the impact of the KL divergence weight during stage-2 RL training. As illustrated in Table 10, removing the KL divergence (w/o KL) results in decreased performance (-3.14%) compared to our full model. Furthermore, we vary the $\beta$ hyperparameter across a range of values [1e-1, 1e-2, 1e-3, 1e-4, 1e-5]. The results indicate that excessively large KL regularization over-constrains the model, while extremely small values provide insufficient constraints. In contrast, an appropriately scaled regularization term, like our 1e-3, offers a balanced trade-off between imposing constraints and encouraging model exploration.
Figure 3: Qualitative comparison. While the Qwen2.5-VL-7B base model and its SFT variant mistakenly grounded the wall and the toilet, respectively, our model correctly understood the instruction and grounded the right object: the shower room.
4.4 Qualitative Visualization
We present a representative case study in Figure 3 to qualitatively evaluate the reasoning and grounding capabilities of our proposed method. The figure demonstrates that while both the Qwen2.5-VL-7B base model and its SFT method fail to accurately ground the target object, our reasoning model successfully accomplishes this task through a coherent two-step reasoning process: (1) identifying the basketball player in Image-1 who requires shower facilities, and (2) precisely localizing the shower room as the appropriate target. In contrast, the base model incorrectly selects the wall, and the SFT method erroneously chooses the toilet. These comparative results provide clear evidence of our model’s enhanced cross-image comprehension and superior instruction-following capabilities.
5 Conclusion
In this paper, we aim to enhance the reasoning and understanding capabilities of MLLMs for real-world multi-image grounding applications. To this end, we propose a post-training strategy incorporating cold-start initialization to establish correct reasoning pathways, complemented by rule-based reinforcement learning to further strengthen the model’s reasoning abilities. Experimental results demonstrate that our approach significantly outperforms both the base model and the baseline SFT method across multiple multi-image grounding and understanding benchmarks. We hope that this work will inspire further research in the multi-image reasoning and grounding community.